A company’s weighted average cost of capital is 12% per year and the market value of its debt is $15 million. The company’s free cash flow last year was $8.5 million and it is expected to grow 20% per year for the next three years. Thereafter, the free cash flow is expected to grow forever at a rate of 5% per year. If the company has two million shares of common stock outstanding, what is the value per share?
Answer #1

Firm value = (8.5*1.2)/1.12 + (8.5*1.2^2)/1.12^2 + (8.5*1.2^3)/1.12^3 + (8.5*1.05)/0.12*1/1.12^3

Firm value = 9.1071428571 + 9.7576530612 + 10.4546282799 + 52.9386559311

Firm value = $82.2580801293 million

Equity value = Firm value - market value of debt

Equity value = 82.2580801293 - 15

Equity value = $67.2580801293 million

Value per share = 67.2580801293/2

Value per share = $33.6290400647

Can you please upvote? Thank You :-)